Without the wheels, a bicycle frame has a mass of 7.77 kg. Each of the wheels can be roughly modeled as a uniform solid disk with a mass of 0.820 kg and a radius of 0.343 m. Find the kinetic energy of the whole bicycle when it is moving forward at 4.05 m/s


1
Expert's answer
2022-03-03T13:22:56-0500

Answer


Total kinetic energy

"KE=\frac{(M+2m)v^2}{2}+2*0.5Iw^2\\=\frac{(777+2*82)*(0.405)^2}{2}+2*0.5*4.82*(1.18)^2\\=83.83J"